The director of research and development is testing a new medicine. She wants to know if there is evidence at the 0.025 level that the medicine relieves pain in less than 335 seconds. For a sample of 15 patients, the mean time in which the medicine relieved pain was 328 seconds with a variance of 576. Assume the population distribution is approximately normal.

State null/alternative hypothesis, test statistic, one/two tailed, decision rule for rejection null hypothesis, reject or fail to reject.
Answer #1

Answer

Here by the problem,

The director of research and development is testing a new medicine. She wants to know if there is evidence at the 0.025 level that the medicine relieves pain in less than 335 seconds. Let us assume \mu be the actual mean pain relieve time of that medicine. We also assume that  the population distribution is approximately normal.

Now for a sample of n=15 patients, the mean time in which the medicine relieved pain was \bar{X}=328 seconds with a variance of s^2=576

So the hypotheses are,

H_0:\mu=335 \ against \ H_a:\mu<335

In order to test that the test statistic be,

t=\frac{\bar{X}-335}{s/\sqrt{n}} where under Ho, t\sim t_{n-1} \ or \ t_{14}

So putting the values we get the test statistic value as,

t_{obs}=\frac{328-335}{24/\sqrt{15}}=-1.1296

Now clearly as the alternative only consider the less than values hence it is a left tailed (one tailed) test.

The p-value of the test statistic be,

p-value=P[t<t_{obs}]=P[t<-1.1296]=0.1388

As the p-value=0.1388>0.025=level of significance,hence based on data we fail to reject null hypothesis and conclude that  the medicine does not relieve pain in less than 335 seconds.

Hence the answer...............

Thank you..............